# DMol: A Schedule-Driven Diffusion Model for Highly Efficient and Versatile Molecule Generation

### v1

@@ 5 DiGress vs DMol
We next compare DiGress and DMol. First, we observe that DiGress also constructs state transition matrices based on marginal distributions, with the key difference with respect to DMol being that its transition matrices vary over time as
@@
where $\overline{\alpha}^{t}=\prod_{\tau=1}^{t}\alpha^{\tau}$ and $\overline{\beta}^{t}=1-\overline{\alpha}^{t}$. Additionally, DiGress also adopts a cosine schedule for $\overline{\alpha}^{t}$$\overline{\alpha}^{t}=\cos^{2}(0.5\pi(t/T+c)/(1+c)),$ where $c>0$ is a constant.
@@
In DMol, for a graph dataset with $n$ nodes, the number of nodes undergoing state changes at time step $t$ equals $N(t)=(1-\alpha)\,n$. At the same time step in Digress, the expected number of nodes undergoing state changes equals $\sum_{i}np^{x}_{i}(1-\alpha)(1-p^{x}_{i})$, where $p^{x}_{i}$ represents the marginal probability of node class $i$. Therefore, at the same time step, the number of nodes modified by DMol is $\frac{1}{\sum_{i}p^{x}_{i}(1-p^{x}_{i})}$ times that expected from Digress (the derivations can be found in Appendix C.2). Moreover, among the nodes whose states are changed by DMol, the proportion of nodes with class $i$ is $p^{x}_{i}$, whereas in DiGress, this proportion equals $p^{x}_{i}(1-p^{x}_{i})$. The ratio of these fractions equals $1-p^{x}_{i}$.
@@
For edges, a similar analysis can be applied, leading to the conclusion that at the last time step $T$, the ratio of the number of edges modified by our noise model and that modified by Digress equals $\frac{r}{\sum_{j}p^{e}_{j}(1-p^{e}_{j})}$, where $r$ is the hyperparameter mentioned in Section 3.2. At each time step, this ratio needs to be adjusted by a correction factor that accounts for the fact that our method only selects edges from the complete graph induced by the chosen nodes. This correction factor is the ratio of the number of edges in the induced subgraph and the total number of edges. Additionally, among the modified edges, the proportion of edges with class index $j$ differs from the same in DiGress by a constant factor of $1-p^{e}_{j}$, where $p^{e}_{j}$ equals the marginal probability of edge class $j$. As a result, DMol can be made to replicate the average “behaviour” of DiGress by simply adjusting specific hyperparameters, whereas the opposite claim is not true. Comparisons of the noise evolution distribution, efficiency and expressivity of DMol and Digress are available in Appendix B and Appendix C.
6 Experiments
We present next the results of running DMol and Compressed DMol [^1] The link to the code: https://github.com/liekon/Discrete-Graph-Generation.git on several benchmarking datasets. For experiments pertaining to other graph models (e.g., SBMs and planar graphs), the reader is referred to Appendix 6.
Benchmarks. We compared DMol and compressed DMol to several graph generation methods. The benchmarking models used in the experiments include GraphRNN (You et al., 2018), GRAN (Liao et al., 2019), GraphVAE (Simonovsky & Komodakis, 2018), GT-VAE (Mitton et al., 2021), Set2GraphVAE (Vignac & Frossard, 2021), SPECTRE (Martinkus et al., 2022b), GraphNVP (Madhawa et al., 2019), GDSS (Jo et al., 2022b), VAE (Gómez-Bombarelli et al., 2018), JT-VAE (Jin et al., 2019), GraphINVENT (Mercado et al., 2021), LSTM (Segler et al., 2018), NAGVAE (Kwon et al., 2020), MCTS (Jensen, 2019) and DiGress (Vignac et al., 2023).
Data. We tested the performance of different generative models on the QM9 (Wu et al., 2018), MOSES (Polykovskiy et al., 2020), and GUACAMOL (Brown et al., 2019) datasets. QM9 is a small dataset that consists of molecules containing up to $9$ heavy atoms; we used $100$K molecules for training, $20$K for validation, and $13$K for testing. Both MOSES and GuacaMol contain millions of molecules, with the number of heavy atoms inside a single molecule bounded by $26$ and $88$, respectively. We use 85% of the molecules for training, 5% for validation, and 10% for testing.
Setup. For each molecule dataset, we run DMol and compressed DMol. For the compressed algorithm, we choose a different (relatively small) number of ring types to be converted into supernodes. The selected ring structures can be found in the Appendix A. We set the hyperparameters of DMol in both setting to $k=2$, $r=0.2$, $\lambda_{1}=5$, $\lambda_{2}=1$, $\lambda_{3}=1$. The number of training epochs is set to $500$. The experiments were conducted using NVIDIA H100 GPUs. The QM9 runs were executed on a single GPU, while the MOSES and GUACAMOL experiments were processed in parallel using $4$ GPUs.
Table 1: Performance comparison on the QM9 dataset.
Model V↑ V.U.↑ V.U.N.↑
GraphVAE 56.1 42.8 26.5
GT-VAE 74.8 16.5 15.6
Set2GraphVAE 60.0 56.8 -
SPECTRE 87.5 31.4 29.0
GraphNVP 83.5 18.4 -
GDSS 96.0 94.6 -
DiGress 97.8 95.2 31.8
CDiGress(Ours) 98.0 95.9 52.8
DMol(Ours) 98.2 95.8 68.6
CDMol(Ours) 98.3 96.0 75.1
Table 2: Performance comparison on MOSES. All graph-based methods except DiGress and DMol have hard-coded rules to ensure high validity. Note that we report slightly lower V values for DiGress compared to the original results, as obtained from our own experiments.
Model V↑ U↑ N↑ Filters↑ FCD↓ SNN↑ Scaf↑
VAE 97.2 99.4 68.6 98.9 0.56 0.62 6.1
JT-VAE 100 100 99.8 96.9 1.08 0.52 9.8
GraphINVENT 96.1 99.5 - 95.8 1.20 0.56 12.4
DiGress 84.8 100 94.5 97.2 1.18 0.55 14.6
CDiGress (Ours) 85.8 100 96.2 97.0 1.16 0.56 14.9
DMol (Ours) 85.6 100 100 97.4 1.15 0.55 14.9
CDMol (Ours) 87.8 100 100 97.8 1.12 0.58 14.8
Table 3: Performance comparison on GUACAMOL.
Model V↑ U↑ N↑ KL div↑ FCD↑
LSTM 96.1 100 91.0 99.3 91.5
NAGVAE 93.2 95.4 100 39.5 1.0
MCTS 100 100 95.2 82.9 1.8
DiGress 84.5 100 99.8 93.1 68.4
CDiGress (Ours) 85.6 100 100 93.0 68.8
DMol (Ours) 85.4 100 100 93.5 68.6
CDMol (Ours) 86.7 100 100 94.2 69.3
Table 4: Performance comparison of molecular metrics on MOSES and GUACAMOL. CL stands for CLscores, ST stands for sample time, and ATRT stands for approximate training time.
Dataset Model CL↑ QED↑ ST(s)↓ ATRT(days)↓
Moses DiGress 4.4965 0.8024 68.24 5.5
CDiGress (Ours) 4.4984 0.8036 75.46 5.5
DMol (Ours) 4.5002 0.8069 5.28 3.0
CDMol (Ours) 4.5033 0.8055 1.52 3.0
Guacamol DiGress 4.0412 0.5650 70.82 6.0
CDiGress (Ours) 4.1902 0.5698 81.36 6.0
DMol (Ours) 4.0425 0.5728 5.49 4.0
CDMol (Ours) 4.2231 0.5786 1.84 4.0
Evaluation Metrics. The metrics used in the experiments include validity, uniqueness, and novelty. Validity (V) measures the proportion of generated molecules that pass basic valency checks (e.g., correspond to valid SMILES files). Uniqueness (U) quantifies the proportion of generated molecules that are distinct. Novelty (N) represents the percentage of generated molecules that do not appear in the training dataset. Furthermore, V.U. stands for the product of validity and uniqueness scores, while V.U.N. stands for the product of validity, uniqueness, and novelty scores.
Since MOSES and GUACAMOL are benchmarking datasets, they introduce their own set of metrics for reporting results. In addition to V, U, and N, MOSES incorporates a filter score, the Frechet ChemNet Distance (FCD) (Preuer et al., 2018), SNN, and scaffold similarity scores, while GUACAMOL uses FCD and KL divergence. These metrics are defined as follows: The filter score measures the proportion of molecules that pass the same filters used to construct the test set. The FCD score measures the similarity between molecules in the training and test sets using embeddings learned by a neural network. SNN quantifies the similarity of a molecule to its nearest neighbor based on the Tanimoto distance. Scaffold similarity compares the frequency distributions of so-called Bemis-Murcko scaffolds. KL divergence measures the differences in the distributions of various physicochemical descriptors.
The metrics mentioned above still have certain limitations. Since these metrics assess the overall quality of the generation process, but do not provide insights into the performance of individual molecules. To address this limitation, we also use the ChEMBL-Likeness Score (CLscore) (Bühlmann & Reymond, 2020) and Quantitative Estimation of Drug-likeness (QED) (Bickerton et al., 2012) score. The CLscore is calculated by considering how many substructures in a molecule also occur in the drug-like dataset ChEMBL (Gaulton et al., 2016). QED evaluates drug-likeness by considering eight widely recognized molecular properties, such as molecular mass and polar surface area. These two scores can be seen as more advanced validity metrics. More details are available in Appendix D.
Other performance metrics we use are the sample time (seconds) needed to generate a batch of $64$ molecules.
Results. The experimental results are presented in Tables 12, and 3, and they pertain to DMol, CDMol (Compressed DMol) proposed in Section 4, and CDiGress (Compressed DiGress). On QM9, DMol achieves the best performance compared to other baseline models. In contrast, the state-of-the-art discrete diffusion model, DiGress, shows a significant deficiency in novelty, indicating that most of the generated molecules are duplicates of those already present in the training set. CDMol retains the V and U scores of DMol, but significantly improves the N score. This is because ring compression makes it more likely for molecules to exhibit diverse structures.
On the MOSES and GUACAMOL datasets, our method achieves a performance comparable to that of the autoregressive model GraphInvent, which leverages chemical software and reinforcement learning for fine-tuning. DMol and CDMol also significantly outperform DiGress (note that the V scores of DiGress on MOSES and Guacamol were reported as $85.7$% and $85.2$%; however, we report the results of our own simulations, for which there is a discrepancy of roughly $1$% against DiGress - such discrepancy is not unexpected, since the output results are random). Although JT-VAE demonstrates excellent V performance, it involves a very complex reconstruction process and potentially suffers from issues associated with VAE-based generative models, such as coarse-graining (which manifests itself as fairly uniform scaffolds of the generated molecules). In contrast, DMol can directly sample molecular graphs with flexible control over the number of generated nodes.
Table 4 presents a comparison between our model and DiGress with respect to the CL score, QED, sample time and approximate training time. The results show that the average CLscore and QED of molecules generated by our model are slightly but uniformly higher than those of DiGress, and that the training and sampling speed are significantly faster. This advantage stems from the fact that our noise model adds noise to subgraphs rather than the entire graph, allowing the denoising model to converge more quickly. Also, our diffusion process uses a number of time steps proportional to the number of nodes in the training graph which further contributes to the reduced execution time. As a parting remark, we point out that although ring compression requires a (simple) decoding step, it significantly reduces the number of sampling steps needed since it effectively makes the compressed molecule graph smaller. However, when ring compression is combined with classical DiGress diffusion, since the total number of sampling steps is fixed, the decoding process actually increases the total sampling time. Therefore, compressed DMol also uniformly outperforms compressed DiGress. Additional results, pertaining to the GruM method (Jo et al., 2024), which were omitted in the main text due to space limitations, along with an ablation study of DMol, are available in Appendix E.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Ring Compression
Since each dataset has a different size of graph and complexity, we choose a different number of rings K to be converted to supernodes. We select $3$ rings for QM9 4, $15$ rings for Moses 5, and $15$ rings for Guacamol 6.
Figure 4: The $3$ selected carbon rings in QM9.
Figure 5: The $15$ selected carbon rings in Moses.
Figure 6: The $15$ selected carbon rings in Guacamol.
Appendix B “Converting” DMol Diffusion into Digress Diffusion
@@ Appendix A Ring Compression
Appendix C Further comparisons between DMol and DiGress
C.1 Noise Distribution Evolution
@@
C.2 Efficiency
To compare the efficiency of DMol and DiGress, we compare the time step increment required to change exactly one node under both settings. In DMol, the number of nodes changing at time $t$ is $N(t)=(1-\alpha^{t})\cdot n$. To find $\delta t_{DMol}$ such that $\delta N(t)=1$, we need
@@ C.2 Efficiency
Following a similar derivation as for DMol, but with a constant $T$ instead of $T=kn$, we arrive at
@@
Assume next, for fairness of comparison, that the maximum number of diffusion step used by DMol and Digress are the same. Then, the ratio of the minimum time steps of interest equals
@@
This proves that $\delta t_{Digress}>\delta t_{DMol}$, or, in words, that DMol requires a smaller time step increment to change exactly one node. As a result, without restrictions on the diffusion times, the number of steps required by DMol to change a node will be smaller than that of required by Digress, which implies that DMol performs graph perturbations more efficiently.
C.3 Expressivity
@@
Appendix D ChEMB Likeness and QED Scores
The ChEMBL-Likeness Score (CLscore) (Bühlmann & Reymond, 2020) is a metric used to assess the drug-likeness of a molecule. It is determined by analyzing how many of the molecule’s substructures are present in the ChEMBL dataset(Gaulton et al., 2016), a database of bioactive molecules with drug-like properties. Essentially, the CLscore quantifies the likelihood of a molecule being drug-like based on its substructure similarity to known compounds in ChEMBL.
Quantitative Estimation of Drug-likeness (QED) (Bickerton et al., 2012) is a computational metric used in drug discovery to assess how “drug-like” a compound is. It combines eight physicochemical properties, including molecular mass ($M_{r}$), octanol–water partition coefficient (ALOGP), number of hydrogen bond donors (HBDs), number of hydrogen bond acceptors (HBAs), molecular polar surface area (PSA), number of rotatable bonds (ROTBs), number of aromatic rings (AROMs), and number of structural alerts (ALERTS), into a single score ranging from 0 (least drug-like) to 1 (most drug-like). The score is measured through defined desirability functions of the eight properties. QED provides a balanced evaluation by weighting these properties based on their distributions in known drugs.
Appendix E Supplementary Experiments
E.1 General Graph Generation
We conducted experiments using the benchmark proposed by (Martinkus et al., 2022a), which comprises two datasets: SBM and Planar. Each dataset consists of 200 graphs. DMol’s ability to accurately model various properties of these graphs was assessed using metrics such as the degree distribution (Deg.), clustering coefficient distribution (Clus.), and orbit count distribution (Orb., the number of occurrences of substructures with 4 nodes), measured by the relative squared Maximum Mean Discrepancy (MMD).

### v3

@@ Appendix B DiGress vs DMol
We now further compare DiGress and DMol. First, we observe that DiGress also constructs state transition matrices based on marginal distributions, with the key difference with respect to DMol being that its transition matrices vary over time as
@@
where $\overline{\alpha}^{t}=\prod_{\tau=1}^{t}\alpha^{\tau}$ and $\overline{\beta}^{t}=1-\overline{\alpha}^{t}$. Additionally, DiGress also adopts a cosine schedule for $\overline{\alpha}^{t}$: $\overline{\alpha}^{t}=\cos^{2}(0.5\pi(t/T+c)/(1+c)),$ where $c>0$ is a constant.
@@
In DMol, for a graph dataset with $n$ nodes, the number of nodes undergoing state changes at time step $t$ equals $N(t)=(1-\alpha)\,n$. At the same time step in Digress, the expected number of nodes undergoing state changes equals $\sum_{i}np^{x}_{i}(1-\alpha)(1-p^{x}_{i})$, where $p^{x}_{i}$ represents the marginal probability of node class $i$. Therefore, at the same time step, the number of nodes modified by DMol is $\frac{1}{\sum_{i}p^{x}_{i}(1-p^{x}_{i})}$ times that expected from Digress (the derivations can be found in Appendix D.2). Moreover, among the nodes whose states are changed by DMol, the proportion of nodes with class $i$ is $p^{x}_{i}$, whereas in DiGress, this proportion equals $p^{x}_{i}(1-p^{x}_{i})$. The ratio of these fractions equals $1-p^{x}_{i}$.
@@
For edges, a similar analysis can be applied, leading to the conclusion that at the last time step $T$, the ratio of the number of edges modified by our noise model and that modified by Digress equals $\frac{r}{\sum_{j}p^{e}_{j}(1-p^{e}_{j})}$, where $r$ is the hyperparameter mentioned in Section 3.2. At each time step, this ratio needs to be adjusted by a correction factor that accounts for the fact that our method only selects edges from the complete graph induced by the chosen nodes. This correction factor is the ratio of the number of edges in the induced subgraph and the total number of edges. Additionally, among the modified edges, the proportion of edges with class index $j$ differs from the same in DiGress by a constant factor of $1-p^{e}_{j}$, where $p^{e}_{j}$ equals the marginal probability of edge class $j$. As a result, DMol can be made to replicate the average “behavior” of DiGress by simply adjusting specific hyperparameters, whereas the opposite claim is not true. Comparisons of the noise evolution distribution, efficiency and expressivity of DMol and Digress are available in Appendix D.
@@
Appendix C Likelihood Computation
DMol can be visualized through the model depicted in Figure  6, informed by the dataset distribution and where nodes are perturbed first, while edges are conditionally perturbed based on the node selection. This framework extends traditional hierarchical variational inference structures by incorporating a crucial dependency relationship: the selection of edges to perturb is conditioned on the selection of nodes at each timestep.
[…]
Figure 6: The conditional noise-addition model of DMol, where $\textbf{K}_{x}(t)$ captures the random node selection while $\textbf{K}_{e}(t)$ captures the random edge selection process; the dashed arrow indicating the dependency of edge selection on node selection.
For a molecular graph $G$, the model likelihood must account for this conditional relationship, according to:
$\log p_{\theta}(G)$ $=\log\sum_{n\in\mathcal{N}}p(n)\int p(G^{T}|n)p_{\theta}(G^{T-1},\ldots,G^{1}|G^{T})p_{\theta}(G|G^{1})d(G^{1},\ldots,G^{T})$ (2)
$=\log p(n_{G})+\log\int p(G^{T}|n_{G})\prod_{t=2}^{T}p_{\theta}(G^{t-1}|G^{t},\textbf{K}_{x}(t),\textbf{K}_{e}(t))p_{\theta}(G|G^{1})d(G^{1},\ldots,G^{T})$ (3)
Note the key difference in the conditional probability $p_{\theta}(G^{t-1}|G_{t},\textbf{K}_{x}(t),\textbf{K}_{e}(t))$, which explicitly accounts for the selection matrices $\textbf{K}_{x}(t)$ and $\textbf{K}_{e}(t)$ that designate which nodes and edges receive perturbations at time $t$.
Following established variational inference principles, we derive a tractable evidence lower bound (ELBO) adapted for this conditional structure:
$\log p_{\theta}(G)\geq\log p(n_{G})$ $+\underbrace{D_{KL}[q(G^{T}|G)\|q_{\mathcal{X}}(n_{G})\times q_{\mathcal{E}}(n_{G})]}_{\text{Prior term}}$ (4)
$+\underbrace{\sum_{t=2}^{T}\mathcal{L}_{t}}_{\text{Diffusion component}}$ (5)
$+\underbrace{E_{q(G^{1}|G)}[\log p_{\theta}(G|G^{1})]}_{\text{Reconstruction component}}$ (6)
Where the diffusion component $\mathcal{L}_{t}$ now explicitly models the conditional relationship between node and edge noise:
$\mathcal{L}_{t}(G)=E_{q(G^{t}|G)}\left[D_{KL}\left[q(G^{t-1}|G^{t},G,\textbf{K}_{x}(t-1),\textbf{K}_{e}(t-1))\|p_{\theta}(G^{t-1}|G^{t},\textbf{K}_{x}(t-1),\textbf{K}_{e}(t-1))\right]\right]$ (7)
This can be further decomposed to highlight the conditional structure:
$\mathcal{L}_{t}=$ $\sum_{i\in K_{x}(t-1)}E_{q(x_{i}^{t}|x_{i})}\left[D_{KL}\left[q(x_{i}^{t-1}|x_{i}^{t},x_{i})\|p_{\theta}(x_{i}^{t-1}|G^{t})\right]\right]$ (8)
$+\sum_{(i,j)\in K_{e}(t-1)}E_{q(e_{ij}^{t}|e_{ij})}\left[D_{KL}\left[q(e_{ij}^{t-1}|e_{ij}^{t},e_{ij})\|p_{\theta}(e_{ij}^{t-1}|G^{t},\textbf{K}_{x}(t-1))\right]\right]$ (9)
Note the critical difference in the edge term, where the predicted probability $p_{\theta}(e_{ij}^{t-1}|G^{t},\textbf{K}_{x}(t-1))$ is explicitly conditioned on the node selection $\textbf{K}_{x}(t-1)$. This formulation captures the key aspect of DMol’s noise process: edges are selected for perturbation only from within the subgraph induced by the selected nodes.
The conditional dependency structure provides two key advantages for DMol: 1. It preserves local structural coherence (e.g, motifs) by ensuring that edge modifications are correlated with node changes. 2. It significantly reduces the computational complexity by restricting edge perturbations to smaller subgraphs.
In practice, each ELBO component remains computationally tractable: we calculate $\log p(n_{G})$ from the empirical node count distributions across the molecular corpus. The prior term involves categorical KL divergence calculations, while the diffusion component now incorporates the conditional selection of edges based on the node selection. The reconstruction component derives from generated probabilities of the original structure given its slightly perturbed version $G^{1}$.
Appendix D Theoretical Analysis of DMol
D.1 Noise Distribution Evolution
@@
D.2 Efficiency
To compare the efficiency of DMol and DiGress, we compare the time step increment required to change exactly one node(minimal noise) under both settings. In DMol, the number of nodes changing at time $t$ is $N(t)=(1-\alpha^{t})\cdot n$. To find $\delta t_{DMol}$ such that $\delta N(t)=1$, we need
@@ D.2 Efficiency
Following a similar derivation as for DMol, we arrive at
@@
Then, the ratio of the minimum time steps of interest equals
@@
This proves that $\delta t_{Digress}>\delta t_{DMol}$, or, in words, that DMol requires a smaller time step increment to change exactly one node. As a result, the number of steps required by DMol to change a node will be smaller than that of required by Digress, which implies that DMol performs graph perturbations more efficiently.
D.3 Expressivity
@@
Appendix E The Metrics used in MOSES and GUACAMOL
The metrics used in MOSES and GUACAMOL dataset are defined as follows: The filter score measures the proportion of molecules that pass the same filters used to construct the test set. The FCD score measures the similarity between molecules in the training and test sets using embeddings learned by a neural network. SNN quantifies the similarity of a molecule to its nearest neighbor based on the Tanimoto distance. Scaffold similarity compares the frequency distributions of so-called Bemis-Murcko scaffolds. KL divergence measures the differences in the distributions of various physicochemical descriptors.
Appendix F ChEMB Likeness, QED Scores, and Shingle Distance
The ChEMBL-Likeness Score (CLscore) Bühlmann and Reymond [2020] is a metric used to assess the drug-likeness of a molecule. It is determined by analyzing how many of the molecule’s substructures are present in the ChEMBL datasetGaulton et al. [2016], a database of bioactive molecules with drug-like properties. Essentially, the CLscore quantifies the likelihood of a molecule being drug-like based on its substructure similarity to known compounds in ChEMBL.
Quantitative Estimation of Drug-likeness (QED) Bickerton et al. [2012] is a computational metric used in drug discovery to assess how “drug-like" a compound is. It combines eight physicochemical properties, including molecular mass ($M_{r}$), octanol–water partition coefficient (ALOGP), number of hydrogen bond donors (HBDs), number of hydrogen bond acceptors (HBAs), molecular polar surface area (PSA), number of rotatable bonds (ROTBs), number of aromatic rings (AROMs), and number of structural alerts (ALERTS), into a single score ranging from 0 (least drug-like) to 1 (most drug-like). The score is measured through defined desirability functions of the eight properties. QED provides a balanced evaluation by weighting these properties based on their distributions in known drugs.
The shingles distance (SD) is a metric designed to quantify the structural divergence between generated molecules and a reference dataset. It is computed by comparing the frequencies of important, predefined molecular drug substructures (motifs), known in RdKit as shingles, according to the ChEMBL dataset Gaulton et al. [2016]. The same set of shingles is used in the computation of the ChEMBL-Likeness Score (CLscore) Bühlmann and Reymond [2020]. SD measures the cosine distance between the vectors of shingle occurrence counts for two molecular distributions. By emphasizing differences in the presence of key drug-relevant motifs, SD offers a more nuanced assessment of both validity and novelty.
Appendix G Supplementary Experiments
G.1 General Graph Generation
We conducted experiments using the benchmark proposed by Martinkus et al. [2022a], which comprises two datasets: SBM and Planar. Each dataset consists of 200 graphs. DMol’s ability to accurately model various properties of these graphs was assessed using metrics such as the degree distribution (Deg.), clustering coefficient distribution (Clus.), and orbit count distribution (Orb., the number of occurrences of substructures with 4 nodes), measured by the relative squared Maximum Mean Discrepancy (MMD).